1.       In a basin area where oil is likely to be found underneath the surface, there are 3 locations with 3 different types of earth compositions, say C1, C2, & C3. The probabilities for these 3 compositions are 0.5, 0.3, & 0.2 respectively. Further, it has been found from the past experience that after drilling of well at these locations, the probabilities of finding oil is 0.2, 0.4, & 0.3 respectively. Suppose, a well is drilled at a location, & it yields oil, what is the probability that the earth composition was C1?



1
Expert's answer
2021-01-25T14:47:24-0500

"P(C_1) = 0.5 \\

P(C_2) = 0.3 \\

P(C_3) = 0.2"

Let E denotes the event of finding the oil. Then

"P(E|C_1) = 0.2 \\

P(E|C_2) = 0.4 \\

P(E|C_3) = 0.3"

Let Ec denotes the event of not finding the oil.


Probability (oil is found):

"P(E) = P(E|C_1)P(C_1) + P(E|C_2)P(C_2)+P(E|C_3)P(C_3) \\

= (0.2 \times 0.5)+(0.4 \times 0.3)+(0.3 \times 0.2) \\

= 0.28"

The probability that the earth composition was C1:

"P(C_1|E) = \frac{P(C_1) \times P(E|C_1)}{P(E)} \\

= \frac{0.5 \times 0.2}{0.28} \\

= 0.357"

Answer: 0.357